# Navigating Distribution Shifts in Medical Image Analysis: A Survey

### v1

@@ Abstract
I Introduction
Medical image analysis (MedIA) [1] has become a cornerstone of modern healthcare, playing a critical role in enhancing diagnostics [2, 3, 4], patient monitoring [5], and treatment planning [6]. With the advent of high-resolution imaging technologies and the increasing complexity of medical data, the application of advanced computational tools has become indispensable. Deep learning (DL) technologies [7, 8, 9, 10], in particular, have revolutionized MedIA by enabling automated and accurate analyses of medical images [11, 12]. These technologies leverage large datasets to train models that can recognize patterns with a precision often surpassing human capabilities [13]. The integration of DL in MedIA not only speeds up diagnostic processes but also offers the potential for personalized healthcare through more accurate patient-specific assessments.
However, the application of deep learning techniques in MedIA faces substantial challenges, primarily due to distribution shifts. These shifts occur because the training data (known as source data) used to develop DL models often come from highly controlled environments or specific populations. When deployed in varied medical settings – like different hospitals, population regions, and time periods – these models encounter data that differ significantly in aspects such as imaging modalities [14], scanning protocols [15], patient populations [16], and temporal changes [17]. These variations expose the models to novel, out-of-distribution patterns (referred to as target data) that they have not been trained to recognize, impairing their ability to generalize effectively; this compromised performance in turn undermines the reliability and effectiveness of DL-based diagnostics. Therefore, addressing these distribution shifts is crucial for the effect and reliable deployment of DL technologies in diverse medical environments.
To this end, this survey focuses on investigating DL-based MedIA under the challenges posed by distribution shifts. In recent years, the research community has actively developed strategies to enhance the adaptability and robustness of DL models. These strategies aim to mitigate the impact of data distribution shifts across diverse medical settings [18, 19]. In real-world healthcare, the successful deployment of DL technologies often encounters various operational constraints that directly leads to different data distribution shift scenarios. These constraints typically stem from several key factors:
@@ I Introduction
    Privacy Concerns: Given the sensitive nature of medical data, privacy concerns [20] revolve around the protection of patient information. These considerations often limit the sharing of medical data among different healthcare institutions, creating data silos that exacerbate the potential data distribution shifts.
@@
    Collaborative Protocols: Collaboration among healthcare institutions enables collective efforts to improve diagnostic models across diverse settings. By adhering to different protocols, various collaborative methods [21, 22] have been developed while meeting specific requirements to alleviate the potential distribution shifts.
@@
Building on these practical considerations while deploying DL models, we categorize existing efforts to manage distribution shifts in MedIA into a hierarchy from simple to hard:
@@
    Joint Training: This approach is feasible when both the source and target data are accessible and there is no privacy concerns. This scenario often occurs when multiple health institutions agree to share their own data, facilitating joint model training [23, 24] and thereby enhancing model adaptability across diverse settings.
@@
    Federated Learning: When multiple institutions seek to cooperate without exposing their distinct datasets due to privacy concerns, federated learning [25] offers a powerful solution. It enables collaborative model improvements across different institutions by training models locally on each dataset and aggregating the learned models without centralizing data storage.
@@
    Fine-tuning: When synchronous collaborations are not allowed for addressing data distribution shifts with privacy concerns, fine-tuning [26, 27] emerges as an effective remedy. This involves using a well pre-trained model and then fine-tuning it on new datasets to transfer learned knowledge to unfamiliar domains.
@@
    Domain Generalization: When data from unseen domains that require model adaptation is inaccessible or unknown, training a model that is generalizable enough to withstand distribution shifts is essential [28, 29]. This involves preparing for unforeseen challenges by developing models that can generalize from the data currently available for training to any potential new environments.
@@
In this survey, we present a nuanced understanding of how deep learning can be strategically deployed to address distribution shifts in MedIA, facilitating the development of diverse and robust applications. While a few surveys have also explored the impact of distribution shifts on MedIA and summarized solutions, our work stands apart in several critical ways. For instance, [18] primarily focuses on domain adaptation (DA) within MedIA, categorizing existing methods based on the degree of DL model supervision while [19] emphasizes Domain Generalization (DG) and organizes existing methods according to common MedIA workflows. Although DA and DG are significant topics with profound impacts on MedIA, these surveys [18, 19] concentrate on the technical aspects of existing approaches, treating MedIA primarily as an application domain. Their classifications, rooted in the intricacies of DL techniques, often overlook the real-world medical constraints that give rise to different distribution shift scenarios. Consequently, they fail to provide a detailed, step-by-step guide addressing the impact of medical data variations.
@@
Unlike them, our approach is grounded in the practical, operational constraints faced by healthcare institutions, examining current DL techniques in light of the real factors affecting MedIA under distribution shifts. Moreover, while some surveys have explored these issues within medical contexts, they often restrict their discussions to specific scenarios (e.g., heart/lung/brain), such as [30, 31, 32, 33], lacking a comprehensive exploration of distribution shifts within MedIA. Our survey addresses these gaps by offering direct and actionable strategies for deploying DL models under the unique operational constraints encountered in real-world applications. This not only serves as a practical guide for medical professionals on employing deep learning to tackle genuine medical challenges but also underscores the transformative potential of DL technologies in MedIA. By highlighting operational constraints and providing tailored solutions, our survey deepens the understanding and broadens the application of deep learning in MeIA, thereby enhancing both the field and the integration of artificial intelligence in healthcare.
TABLE I: Definition of mathematical notations.
Notation Definition Notation Definition
@@
$\mathcal{X},\mathcal{Z},\mathcal{Y}$ Input, feature, output space $P(\cdot)$ Prior Knowledge
@@
$\mathbf{x},\mathbf{z},y$ Input, feature, label variables $\mathcal{L}(\cdot,\cdot)$ Loss function
@@
$\mathcal{D}_{t}$ Target Domain $q_{\theta}(\cdot)$ Predictive function
$\mathcal{D}_{s}$ Source Domain $\mathcal{M}(\cdot)$ Manipulation function
$p(\cdot)$ Probability distribution $f(\cdot)$ Feature mapping function
I-A Problem Definition
@@
In this section, we formalize the distribution shift problem with notations defined in Table I for easy reading. A domain $\mathcal{D}$ is a joint distribution $p(x,y)$ defined on the input-output space $\mathcal{X}\times\mathcal{Y}$, where random variables $x\in\mathcal{X}$ and $y\in\mathcal{Y}$ denote the input data and the output label, respectively. We typically deal with two distinct datasets, known as the source and target domains. The Source Domain $\mathcal{D}_{s}=\{(x,y)\sim p_{s}(x,y)\}$. comprises medical images $x$ such as X-rays or MRI scans, each paired with a label $y$ that might be categorical information regarding disease diagnosis or the segmentation mask. The Target Domain $\mathcal{D}_{t}=\{(x,y)\sim p_{t}(x,y)\}$ originates from a different but related distribution to that of the source. For instance, they might come from different medical imaging devices or patient populations. Note that for both source and target distributions, $p_{s}(x,y)=p_{s}(x)p_{s}(y|x)$ and $p_{t}(x,y)=p_{t}(x)p_{t}(y|x)$. We take the standard covariate shift assumption as Distribution Shift, i.e., $p_{s}(y|x)=p_{t}(y|x)$ and $p_{s}(x)\neq p_{t}(x)$. In this situation, the model $q_{{\theta}}(y|x)$ solely trained on the source domain cannot well represent the true, domain-invariant distribution $p(y|x)$. Therefore, a variety of research concentrates on adjusting $q_{{\theta}}(y|x)$ to maximize its predictive performance on the target distribution.
II Background
II-A Distribution Shifts in Medical Image Analysis
@@
The efficacy of DL models largely hinges on the assumption that the training and testing data are independently and identically distributed (i.i.d). However, this assumption often does not hold in the complex and diverse environment of clinical practice. The inherent heterogeneity in medical imaging, arising from different modalities, varying protocols, diverse patient demographics, and temporal shifts, introduces significant distribution shifts. In the following, we provide a concise illustration (see visual examples in Fig. 1):
@@
    Temporal Shifts: Longitudinal studies and data collected over extended periods often encounter shifts due to the progression of diseases, the impact of treatments, and changes in physiological states. As a result, models trained on historical data may not perform optimally on current or future data. These shifts challenge the performance of models when applied outside their training domain, a common occurrence given the variety of imaging techniques used in healthcare.
TABLE II: Taxonomy of DL techniques designed for MedIA under Distribution Shifts.
@@
III Categorization and Frameworks
This section provides an overview of our categorization rationale and framework for addressing MedIA under distribution shifts. We primarily consider the real-world operational constraints encountered when deploying DL techniques in MedIA, such as Data Accessibility, Privacy Concerns, and Collaborative Protocols. These factors shape different scenarios under distribution shift in MedIA, leading to the classification of exisiting DL techniques into four main categories: Joint Training, Federated Learning, Fine-tuning, and Domain Generalization. A visual depiction of this classification is provided in Fig. 2. Within each major category, we further subdivide the techniques into prominent subdomains in the field, primarily based on differences in label rate and diversity of data. These subdomains are then ranked according to their learning difficulty (see Table II). Furthermore, to help researchers quickly identify the technical types of methods they are interested in, we categorize the methods within each subdomain under three aspects as elaborated in the following:
@@ III Categorization and Frameworks
It is important to note that although methods are classified by their primary innovation, many integrate multiple strategies across these aspects. This structured taxonomy not only aids researchers in identifying methods suited to specific problems but also facilitates a comparative and systematic analysis.
Figure 1: Illustration of medical imaging distribution shifts, showcasing from Imaging Modalities (Cardiac Substructure [14]), Scanning Protocols (Cross-Site Prostate [15]), Patient Demographics (Cross-Population Chest [34, 35]), and Temporal Shifts (Mouse Lung [36]).
Figure 2: The diagram categorizes existing deep learning techniques into four main approaches, each addressing real-world operational constraints including Data Accessibility, Privacy Concerns, and Collaborative Protocols. In Joint Training, hospitals collaborate by sharing data for model training. Federated Learning enables collaboration without direct data sharing, maintaining privacy. Fine-tuning adapts pre-trained models from one institution to another. Domain Generalization develops models that generalize across diverse settings, even without access to target data, to mitigate distribution shifts.
IV Joint Training
Joint Training is a crucial domain adaptation strategy in MedIA, particularly effective when target data is freely accessible and privacy concerns are minimal. This method excels in environments where healthcare institutions can collaboratively share data, creating the ideal conditions for joint model training. Such collaboration significantly enhances the adaptability of models across varied medical settings by integrating both source and target data. Typically, the source dataset is fully-labeled, whereas the target dataset often exhibits varying labeling rates due to changes in medical scenarios, introducing complexities to DL model training. In response, a variety of joint training strategies have emerged, each designed to address the specific challenges posed by fluctuating label availability on target data. These methods are categorized based on the level of target supervision, ranging from Supervised to Semi-supervised, and Unsupervised Joint Training.
IV-A Supervised Joint Training
@@
IV-A1 Data Management
Cross-modal Translation. Cross-modal translation plays a pivotal role in addressing the challenge of integrating data from diverse imaging modalities, which often exhibit distinct intensity and texture characteristics. This technique facilitates the conversion of data between modalities, such as from MRI to CT images, enabling the use of a unified dataset for training despite the inherent discrepancies. By synthesizing data from one modality in the form that resembles another, cross-modal translation helps to overcome the shortage data problem and enhances the robustness of the training process. Specifically, Generative Adversarial Networks have proven to be particularly effective for cross-modal translation [37, 38, 39, 40] by creating high-quality synthetic images that maintain the domain-specific characteristics of the target modality. For example, the shape-consistency approach [37] leverages GANs for volume-to-volume translation, ensuring that the structural integrity of medical images is preserved across modalities.
IV-A2 Model Design
Architecture Variations. Novel architectural designs are crucial for addressing domain adaptation challenges. For instance, the domain-adaptive two-stream U-Net, applied for electron microscopy image segmentation [41], features a dual-stream architecture that supports selective weight sharing between source and target domains. This design enhances adaptability by allowing the model to fine-tune its responses to the unique characteristics of each domain. Similarly, the Multi-Site Network (MS-Net) for cross-site prostate segmentation [23] incorporates Domain-Specific Batch Normalization (DSBN). DSBN effectively manages inter-site variability by providing distinct feature normalization for each site, ensuring that the model remains robust across diverse MRI datasets.
IV-A3 Optimization Strategy
Metric Learning. Metric Learning has proven instrumental in maintaining high generalization performance across different data domains. A notable application is demonstrated in [42], where metric learning is employed to enhance domain adaptation for Wireless Capsule Endoscopy. This approach utilizes triplet loss, a form of contrastive learning, which effectively minimizes the distance between embeddings of samples with the same labels from different domains while maximizing the distance between samples with different labels from the same domains. By doing so, it ensures that the model can accurately interpret and classify medical images regardless of the specific device version.
IV-B Semi-supervised Joint Training

### v2

@@ Abstract.
1. Introduction
Medical image analysis (MedIA) (Duncan and Ayache, 2000; Li et al., 2023d) has become a cornerstone of modern healthcare, playing a critical role in enhancing diagnostics (Liu et al., 2022d; Jasti et al., 2022), patient monitoring (Gozes et al., 2020), and treatment planning (Jin et al., 2021). With the advent of high-resolution imaging technologies and the increasing complexity of medical data, the application of advanced computational tools has become indispensable. Deep learning (DL) technologies (Huang et al., 2019; Ronneberger et al., 2015; Manzari et al., 2023; Ma et al., 2024), in particular, have revolutionized MedIA by enabling automated and accurate analyses of medical images (Shen et al., 2017; Anwar et al., 2018; Chen et al., 2021b; Carrasco et al., 2024; Hertel and Benlamri, 2023; Kumar et al., 2018; Avola et al., 2021; Guo et al., 2016). These technologies leverage large datasets to train models that can recognize patterns with a precision often surpassing human capabilities (Malik et al., 2020). The integration of DL in MedIA not only speeds up diagnostic processes but also offers the potential for personalized healthcare through more accurate patient-specific assessments.
However, the application of deep learning techniques in MedIA faces substantial challenges, primarily due to distributicon shifts. These shifts occur because the training data (known as source data) used to develop DL models often come from highly controlled environments or specific populations. When deployed in varied medical settings – like different hospitals, population regions, and time periods – these models encounter data that differ significantly in aspects such as imaging modalities (Zhuang and Shen, 2016), scanning protocols (Liu et al., 2020a), patient populations (Fu et al., 2024), and temporal changes (Yang et al., 2019b). These variations expose the models to novel, out-of-distribution patterns (referred to as target data) that they have not been trained to recognize, impairing their ability to generalize effectively; this compromised performance in turn undermines the reliability and effectiveness of DL-based diagnostics. Therefore, addressing these distribution shifts is crucial for the effect and reliable deployment of DL technologies in diverse medical environments.
To this end, this survey focuses on investigating DL-based MedIA under the challenges posed by distribution shifts. In recent years, the research community has actively developed strategies to enhance the adaptability and robustness of DL models. These strategies aim to mitigate the impact of data distribution shifts across diverse medical settings (Guan and Liu, 2021; Yoon et al., 2023). In real-world healthcare, the successful deployment of DL technologies often encounters various operational constraints that directly leads to different data distribution shift scenarios. These constraints typically stem from several key factors:
@@ 1. Introduction
    Privacy Concerns: Given the sensitive nature of medical data, privacy concerns (Kaissis et al., 2020) revolve around the protection of patient information. These considerations often limit the sharing of medical data among different healthcare institutions, creating data silos that exacerbate the potential data distribution shifts.
@@
    Collaborative Protocols: Collaboration among healthcare institutions enables collective efforts to improve diagnostic models across diverse settings. By adhering to different protocols, various collaborative methods (Lenga et al., 2020a; Ng et al., 2021) have been developed while meeting specific requirements to alleviate the potential distribution shifts.
Figure 1. The diagram categorizes existing deep learning techniques into four main approaches, each addressing real-world operational constraints including Data Accessibility, Privacy Concerns, and Collaborative Protocols. In Joint Training, hospitals collaborate by sharing data for model training. Federated Learning enables collaboration without direct data sharing, maintaining privacy. Fine-tuning adapts pre-trained models from one institution to another. Domain Generalization develops models that generalize across diverse settings, even without access to target data, to mitigate distribution shifts.
@@
Building on these practical considerations while deploying DL models, we categorize existing efforts to manage distribution shifts in MedIA into a hierarchy from simple to hard (see Fig. 1):
@@
    Joint Training: This approach is feasible when both the source and target data are accessible and there is no privacy concerns. This scenario often occurs when multiple health institutions agree to share their own data, facilitating joint model training (Liu et al., 2020b; Chen et al., 2019a) and thereby enhancing model adaptability across diverse settings.
@@
    Federated Learning: When multiple institutions seek to cooperate without exposing their distinct datasets due to privacy concerns, federated learning (Nguyen et al., 2022; McMahan et al., 2017) offers a powerful solution. It enables collaborative model improvements across different institutions by training models locally and aggregating the learned models without centralizing data storage.
@@
    Fine-tuning: When synchronous collaborations are not allowed for addressing data distribution shifts with privacy concerns, fine-tuning (Hosseini-Asl et al., 2016; Yang et al., 2022a) emerges as an effective remedy. This involves using a well pre-trained model and then fine-tuning it on new datasets to transfer learned knowledge to unfamiliar domains.
@@
    Domain Generalization: When data from unseen domains that require model adaptation is inaccessible or unknown, training a model that is generalizable enough to withstand distribution shifts is essential (Li et al., 2022a; Zhou et al., 2022a). This involves preparing for unforeseen challenges by developing models that can generalize from the data currently available for training to any potential new environments.
@@
In this survey, we present a nuanced understanding of how deep learning can be strategically deployed to address distribution shifts in MedIA, facilitating the development of diverse and robust applications. While a few surveys have also explored the impact of distribution shifts on MedIA and summarized solutions, our work stands apart in several critical ways. For instance, (Guan and Liu, 2021) primarily focuses on domain adaptation (DA) within MedIA, categorizing existing methods based on the degree of DL model supervision while (Yoon et al., 2023) emphasizes Domain Generalization (DG) and organizes existing methods according to common MedIA workflows. Although DA and DG are significant topics with profound impacts on MedIA, these surveys (Guan and Liu, 2021; Yoon et al., 2023) concentrate on the technical aspects of existing approaches, treating MedIA primarily as an application domain. Their classifications, rooted in the intricacies of DL techniques, often overlook the real-world medical constraints that give rise to different distribution shift scenarios. Consequently, they fail to provide a detailed, step-by-step guide addressing the impact of medical data variations.
@@
Unlike them, our approach is grounded in the practical, operational constraints faced by healthcare institutions, examining current DL techniques in light of the real factors affecting MedIA under distribution shifts. Moreover, while some surveys have explored these issues within medical contexts, they often restrict their discussions to specific scenarios (e.g., heart/lung/brain), such as (Xie et al., 2023; Li et al., 2023a; Kora et al., 2022; Yu et al., 2022; Kim et al., 2022), lacking a comprehensive exploration of distribution shifts within MedIA. Our survey addresses these gaps by offering direct and actionable strategies for deploying DL models under the unique operational constraints encountered in real-world applications. This not only serves as a practical guide for medical professionals on employing deep learning to tackle genuine medical challenges but also underscores the transformative potential of DL technologies in MedIA. By highlighting operational constraints and providing tailored solutions, our survey deepens the understanding and broadens the application of deep learning in MeIA, thereby enhancing both the field and the integration of artificial intelligence in healthcare.
2. Preliminaries
Table 1. Definition of mathematical notations.
Notation Definition Notation Definition Notation Definition
@@ 2. Preliminaries
$\mathcal{X},\mathcal{Z},\mathcal{Y}$ Input, feature, output space $p(\cdot)$ Probability distribution $\mathcal{M}(\cdot)$ Manipulation function
@@
$\mathbf{x},\mathbf{z},y$ Input, feature, label variables $P(\cdot)$ Prior Knowledge $f(\cdot)$ Feature mapping function
@@
$\mathcal{D}_{t},\mathcal{D}_{s}$ Target, Source Domains $q_{\theta}(\cdot)$ Predictive function $\mathcal{L}(\cdot,\cdot)$ Loss function
@@
In this section, we formalize the distribution shift problem with notations defined in Table 1 for easily reading. A domain $\mathcal{D}$ is a joint distribution $p(x,y)$ defined on the input-output space $\mathcal{X}\times\mathcal{Y}$, where random variables $x\in\mathcal{X}$ and $y\in\mathcal{Y}$ denote the input data and the output label, respectively. We typically deal with two distinct datasets, known as the source and target domains. The Source Domain $\mathcal{D}_{s}=\{(x,y)\sim p_{s}(x,y)\}$. comprises medical images $x$ such as X-rays or MRI scans, each paired with a label $y$ that might be categorical information regarding disease diagnosis or the segmentation mask. The Target Domain $\mathcal{D}_{t}=\{(x,y)\sim p_{t}(x,y)\}$ originates from a different but related distribution to that of the source. For instance, they might come from different medical imaging devices or patient populations. Note that for both source and target distributions, $p_{s}(x,y)=p_{s}(x)p_{s}(y|x)$ and $p_{t}(x,y)=p_{t}(x)p_{t}(y|x)$. We take the standard covariate shift assumption as Distribution Shift, i.e., $p_{s}(y|x)=p_{t}(y|x)$ and $p_{s}(x)\neq p_{t}(x)$. In this situation, the model $q_{{\theta}}(y|x)$ solely trained on the source domain cannot well represent the true, domain-invariant distribution $p(y|x)$. Therefore, a variety of research concentrates on adjusting $q_{{\theta}}(y|x)$ to maximize its predictive performance on the target distribution.
3. Background
Figure 2. Illustration of medical imaging distribution shifts, showcasing from Imaging Modalities (Cardiac Substructure (Zhuang and Shen, 2016)), Scanning Protocols (Cross-Site Prostate (Liu et al., 2020a)), Patient Demographics (Cross-Population Chest (Kermany et al., 2018; Wang et al., 2017)), and Temporal Shifts (Mouse Lung (Namati et al., 2010)).
3.1. Distribution Shifts in Medical Image Analysis
@@
The efficacy of DL models largely hinges on the assumption that the training and testing data are independently and identically distributed (i.i.d). However, this assumption often does not hold in the complex and diverse environment of clinical practice. The inherent heterogeneity in medical imaging, arising from different modalities, varying protocols, diverse patient demographics, and temporal shifts, introduces significant distribution shifts. In the following, we provide a concise illustration with visual examples presented in Fig. 2:
@@
    Temporal Shifts: Longitudinal data collected over extended periods often encounter shifts due to physiological states change, treatment impact, and the progression of diseases. As a result, models trained on historical data may not perform optimally on current or future data.
Table 2. Taxonomy of DL techniques designed for MedIA under Distribution Shifts.
@@
4. Categorization and Frameworks
This section provides an overview of our categorization rationale and framework for addressing MedIA under distribution shifts. We primarily consider the real-world operational constraints encountered when deploying DL techniques in MedIA, such as Data Accessibility, Privacy Concerns, and Collaborative Protocols. These factors shape different scenarios under distribution shift in MedIA, leading to the classification of exisiting DL techniques into four main categories: Joint Training, Federated Learning, Fine-tuning, and Domain Generalization (see Table 2). Within each major category, we further subdivide the techniques into prominent subdomains in the field, primarily based on differences in label rate and diversity of data. These subdomains are then ranked according to their learning difficulty. Furthermore, to help researchers quickly identify the technical types of methods they are interested in, we categorize the methods within each subdomain under three aspects:
@@ 4. Categorization and Frameworks
It is important to note that while methods are classified by their primary innovation, many integrate multiple strategies across these aspects. This structured taxonomy not only aids in identifying methods suited to specific problems but also facilitates a comparative and systematic analysis.
5. Joint Training
Joint Training is a crucial domain adaptation strategy in MedIA, particularly effective when target data is freely accessible and privacy concerns are minimal. This method excels in environments where healthcare institutions can collaboratively share data, creating the ideal conditions for joint model training. Such collaboration significantly enhances the adaptability of models across varied medical settings by integrating both source and target data, as shown in Fig. 3. Typically, the source dataset is fully-labeled, whereas the target dataset often exhibits varying labeling rates due to changes in medical scenarios, introducing complexities to DL model training. In response, a variety of joint training strategies have emerged, each designed to address the specific challenges posed by fluctuating label availability on target data. These methods are categorized based on the level of target supervision, ranging from Supervised to Semi-supervised, and Unsupervised Joint Training.
5.1. Supervised Joint Training
@@
5.1.1. Data Management
5.1.1.1 Cross-modal Translation.
Cross-modal translation plays a pivotal role in addressing the challenge of integrating data from diverse imaging modalities, which often exhibit distinct intensity and texture characteristics. This technique facilitates the conversion of data between modalities, such as from MRI to CT images, enabling the use of a unified dataset for training despite the inherent discrepancies. By synthesizing data from one modality in the form that resembles another, cross-modal translation helps to overcome the shortage data problem and enhances the robustness of the training process. Specifically, Generative Adversarial Networks have proven to be particularly effective for cross-modal translation (Cai et al., 2019; Li et al., 2020f; Zhu et al., 2021; Tomar et al., 2022a) by creating high-quality synthetic images that maintain the domain-specific characteristics of the target modality. For example, the shape-consistency approach (Cai et al., 2019) leverages GANs for volume-to-volume translation, ensuring that the structural integrity of medical images is preserved across modalities.
5.1.2. Model Design
5.1.2.1 Architecture Variations.
Novel architectural designs are crucial for addressing domain adaptation challenges. For instance, the domain-adaptive two-stream U-Net, applied for electron microscopy image segmentation (Bermúdez-Chacón et al., 2018), features a dual-stream architecture that supports selective weight sharing between source and target domains. This design enhances adaptability by allowing the model to fine-tune its responses to the unique characteristics of each domain. Similarly, the Multi-Site Network (MS-Net) for cross-site prostate segmentation (Liu et al., 2020b) incorporates Domain-Specific Batch Normalization (DSBN). DSBN effectively manages inter-site variability by providing distinct feature normalization for each site, ensuring that the model remains robust across diverse MRI datasets.
5.1.3. Optimization Strategy
5.1.3.1 Metric Learning.
Metric Learning has proven instrumental in maintaining high generalization performance across different data domains. A notable application is demonstrated in (Laiz et al., 2019), where metric learning is employed to enhance domain adaptation for Wireless Capsule Endoscopy. This approach utilizes triplet loss, a form of contrastive learning, which effectively minimizes the distance between embeddings of samples with the same labels from different domains while maximizing the distance between samples with different labels from the same domains. By doing so, it ensures that the model can accurately interpret and classify medical images regardless of the specific device version.
Figure 3. Joint Training in MedIA: It enables data sharing among healthcare institutions without privacy constraints, integrating source and target data to improve model adaptability across diverse medical settings. Images are collected from datasets Kather16 and Kather19 (Kather et al., 2016, 2019).
5.2. Semi-supervised Joint Training